It is known that quantity demanded decreases by two units for each $1 increase in price. At a price of $5, quantity demanded is ten units. With the solution)
a) What will be the quantity demanded is price is zero?
b) Write and equation for quantity demanded as a function of price.
c) Write an equation that expresses price as a function of quantity.
d) Write an equation for total revenue.
1
Expert's answer
2020-08-02T15:54:08-0400

It is known that quantity demanded decreases by two units for each $1 increase in price. At a price of $5, quantity demanded is ten units.

a) If the price is zero the quantity demanded will be zero too.

b) The equation for quantity demanded as a function of price is "Q = 20 - 2P."

c) An equation that expresses price as a function of quantity will be "P = 10 - 0.5Q."

d) An equation for total revenue is: "TR = P*Q = (10 - 0.5Q)*Q = 10Q - 0.5Q^2."